A 2.00 × 102 kg iceboat is propelled across the horizontal surface of a frozen lake by the wind. The wind exerts a constant force of 4.00 × 102 N while the boat moves 0.90 km. Assume that frictional forces are negligible and that the boat starts from rest. What is the boat’s final speed?
1
Expert's answer
2013-01-14T10:53:45-0500
The work done by a constant force of magnitude F ona point that moves a distance d in the direction of the forceis the product,
A = F*d
The law of conservationof energy:
A = T
T - kinetic energy
T = m*v^2/2
F*d = m*v^2/2
v = Sqrt[2*F*d/m] =Sqrt[2*4*102 H *900 m/(2*102 kg) ] = Sqrt[ 4*900 ] = 60 m/s